OpenStudy (eric_d):

Differentiate following with respect to x a) ln ( x^2 + 1 )

OpenStudy (eric_d):

@Callisto

OpenStudy (callisto):

Chain rule.

OpenStudy (callisto):

u = x^2 +1

OpenStudy (jhannybean):

\[d(\ln(u)) = \frac{1}{u} \cdot u'\]

OpenStudy (eric_d):

 like this ?